1. Name the property of real numbers illustrated below. Be specific with the naming (include "of addition" or "of multiplication," if applicable).
(-8)(13) = (13)(-8)



2. Combine like terms:
-3x2 + 2x – 4x2 – 9 + 6x – 2x2 + 8


3. Solve for x:
8x – 11 = -11x + 18


4. Solve for x:
-2(x – 5) + 7 = x – 8 – 5x

Answer by stanbon(75887)

1. Name the property of real numbers illustrated below. Be specific with the naming (include "of addition" or "of multiplication," if applicable).
(-8)(13) = (13)(-8)
commutative; multiplication
--------------------------

2. Combine like terms:
-3x2 + 2x – 4x2 – 9 + 6x – 2x2 + 8
-9x^2+8x-1
-------------
3. Solve for x:
8x – 11 = -11x + 18
19x = 29
x = 29/19
-------------------

4. Solve for x:
-2(x – 5) + 7 = x – 8 – 5x
-2x+10+7 = -4x-8
2x = -25
x = -12 1/2
================
Cheers,
Stan H.